OpenStudy (anonymous):

u(x) = x^4 – 5.05x^3 – 8.295x^2 + 34.812x – 23.1768

OpenStudy (saifoo.khan):

what u wanna do? o.O

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

the polynomial has to be factored

OpenStudy (anonymous):

Rationalize the fractions:\[x^4 -\frac{101}{20}*x^3-\frac{1659}{200}*x^2 +\frac{8703}{250}*x - \frac{28971}{1250} \]Factor the expression.\[\frac{(5 x-6)^2 (10 x+29) (20 x-111)}{5000} \]\[x=\frac{6}{5},\frac{6}{5},\frac{-29}{10},\frac{111}{20} \]

OpenStudy (anonymous):

thanks robtobey your a lifesaver

OpenStudy (anonymous):

question how did u do it?

OpenStudy (anonymous):

Used Mathematica to rationalize the decimal fractions with the Rationalize function. For example, Rationalize[23.1768] =28971/1250 When the expression was finished the expression was highlighted with a mouse and a click on the "Factor" button was executed. The numerator over 5000 appeared in a small fraction of a second. I was 77 years old 3 days ago and I need all the help I can get.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Thank you very much.

OpenStudy (anonymous):

Your Welcome.